# CropCraft: Inverse Procedural Modeling for 3D Reconstruction of Crop Plants

### v1

@@ 4.1 Canopy Reconstruction
The Poisson reconstruction fails because it completely ignores occluded regions and also frequently stitches nearby leaves together, creating extraneous faces. The MLP gives reasonable predictions in terms of soybean leaf area, but tends to give highly erroneous leaf angles as a result of the domain gap between the synthetic histograms it was trained on and the actual histograms from the NeRF depth maps. This is exacerbated with the maize data, where the NeRF renderings are blurrier. The trust-region optimization produces leaf angles comparable with our method, but does not perform as well for leaf area, most likely due to falling into local minima or making erroneous steps due to the stochasticity of the procedural generation model.
@@ 4.2 Ablation Studies
✓ | ✓ | ✓ | ✓ | 0.69 | 0.15 | 12.07 | 7.39 |
We conduct an ablation study on the components of our loss function in Tab. 2. We observe that each of the components contribute to producing a more accurate reconstruction, some significantly more than others. The depth derivative loss has a pronounced beneficial effect on the leaf area metrics. This is because the other statistics may have trouble constraining the leaf size when the canopy has covered most of the ground, while the depth derivative distribution will be strongly shifted upwards with smaller leaves compared to larger ones. Interestingly, removing most of the other terms one at a time does not actually significantly harm the final performance. The hard constraints imposed by the procedural model and the soft constraints imposed by the remaining terms are able to force a similar solution. We still decide to include all the terms in our final method due to the improved LAIPE, which we prefer over LAIE as it is not biased towards the later time points with larger plants.
@@ 4.3 Photosynthesis Simulation
The outputs of our method can be directly used with radiative transfer models to predict photosynthesis rates, which directly impacts crop productivity. Using Helios [3], a state-of-the-art biophysical modeling framework, we perform radiative transfer simulations on the soybean canopies reconstructed by our method. Helios can predict the net photosynthesis rate for every leaf mesh face at any point in time given a set of environmental variables measured by a flux tower with sensors for temperature, humidity, radiation, etc. We show visualizations of the per-face photosynthesis rate as well as timeseries of the aggregate rate across the entire canopy in Fig. 6. Ground-truth values for the photosynthesis rate are usually calculated at a landscape-level using eddy covariance data from a flux tower [32]. Since we do not have paired flux tower measurements for our dataset (collected in 2023), we use flux tower measurements from a soybean field in the previous year (2022). Because the measurements were collected in a similar location and climate as our data, we can reasonably expect that they reflect the corresponding variables for the field in our data.
The results suggest that the predictions made by the model can accurately represent the net photosynthesis over time as well as its spatial variations throughout the canopy. This validates the potential of our reconstruction pipeline for facilitating large-scale monitoring of crop productivity.
4.4 Limitations
One limitation of our method is that the inverse procedural modeling is dependent on the NeRF reconstruction performance, which may degrade due to wind-induced leaf motion. Another limitation is that the procedural generation models used cannot model certain details, e.g. damaged leaves and non-leaf organs of the plants. A final limitation is that our RANSAC-based row-fitting method is sensitive to hyperparameters such as the inlier threshold and segmentation threshold, which may need to be tuned per dataset.
@@ 5 Conclusion
We presented a novel method for 3D reconstruction of agricultural crops that combines conventional multi-view reconstruction with inverse procedural modeling to produce complete plant shapes despite heavy occlusion. Our method first constructs a neural radiance field (NeRF) and then optimizes a procedurally generated model to be consistent with the NeRF in a visibility-aware manner. We validate our method on data collected in real-world agricultural fields and show that it can reconstruct realistic crop canopies across a variety of plant growth stages. We also show that the reconstructions can enable realistic simulations of crop photosynthesis using a radiative transfer model. The method could potentially be improved in the future by incorporating plant growth priors for temporal consistency, or optimizing the model in a coarse-to-fine manner to enable reconstructing more shape details.
@@ Acknowledgments.
This work is supported by NSF Awards #2331878, #2340254, #2312102, #2414227, and #2404385, as well as an Intel research gift, the NCSA Faculty Fellowship, and the Agroecosystem Sustainability Center at UIUC.
@@
A Additional Experimental Details
A.1 Additional Procedural Model Details
@@
Our procedural soybean model is based on the mCanopy model from Song et al. [42]. Each plant consists of a main stem with up to 14 (trifoliate) nodes, and up to 6 branch stems with up to 2 nodes each. Each node consists of three coplanar leaves, with the middle leaf connected by a small petiole, and is connected to its stem by a larger petiole. The leaf length (and width), the petiole length, the angle with the stem, and the distance between nodes are variable, as they are affected by the multiplier parameters. On the main stem, when all the multipliers are set to 1, the variables follow a profile from top to bottom based on field data measured by Song et al. [42]. Branches start growing with every new main stem node starting from the 8th, with one branch per main stem node starting from the bottom and up to the 6th. The number of nodes per branch is selected randomly according to the distribution defined by Song et al. [42]. The variables for the branch nodes are the same across nodes (but are affected by the multipliers). When the number of nodes is not an integer, it is rounded up and the last node is scaled down in size linearly according to the remainder. Additionally, random noise is added to each petiole angle following a Gaussian distribution with mean 0 and standard deviation 5 degrees. Each node grows on alternating sides of their stem, with additional azimuthal rotation following a Gaussian distribution with mean 0 and standard deviation 60 degrees. The range of possible values for each model parameter are provided in Tab 3.
@@
A.2 Additional Row Fitting and Depth Rendering Details
@@
A.3 Mask and Histogram Details
@@
A.4 Baseline Implementation Details
@@
B Additional Results
B.1 Additional Qualitative Results
@@
B.2 Bayesian Optimization Performance Analysis

### v2

@@ 4.1 Canopy Shape Modeling
The Poisson reconstruction fails because it completely ignores occluded regions and also frequently stitches nearby leaves together, creating extraneous faces. The MLP gives reasonable predictions in terms of soybean leaf area, but tends to give highly erroneous leaf angles as a result of the domain gap between the synthetic histograms it was trained on and the actual histograms from the NeRF depth maps. This is exacerbated with the maize data, where the NeRF renderings are blurrier. The trust-region optimization produces leaf angles comparable with our (Bayesian optimization) method, but does not perform as well for leaf area, most likely falling into local minima or making erroneous steps due to the stochasticity of the procedural model.
Naive geometric optimization.
We emphasize that our goal is not exact leaf-level reconstruction of the crop canopy. Instead, our task demands complete crop canopies that capture key shape statistics of the actual leaf distribution, such that downstream applications like growth monitoring or photosynthesis prediction can be conducted accurately. Our model parameterizations and loss function terms are designed to capture precisely these key shape variations, and not overfit to irrelevant geometric details of the input observations. To illustrate this point, we conduct an experiment using differentiable rendering to directly optimize the vertices of our output mesh to further fit the observed depth map. The optimization is done in PyTorch3D [43] using SGD with learning rate 0.01 and momentum 0.9 on an MSE loss. The results are shown in Fig. 6. Although the pixel-wise depth error can be made very low, important variables such as the total leaf area become less accurate, which is reflected in the LAI [39]. This justifies the need for shape constraints and distribution-level optimization.
@@ 4.2 Ablation Studies
[…]
We conduct an ablation study on the components of our loss function in Tab. 2. We observe that each of the components contribute to producing a more accurate reconstruction, some significantly more than others. The depth derivative loss has a pronounced beneficial effect on the leaf area metrics. This is because the other statistics may have trouble constraining the leaf size when the canopy has covered most of the ground, while the depth derivative distribution will be strongly shifted upwards with smaller leaves compared to larger ones. We note that removing some of the other terms one at a time does not actually significantly harm the final performance. The hard constraints imposed by the procedural model and the soft constraints imposed by the remaining terms are able to force a similar solution. We decide to include all the terms in our final method due to the improved LAIPE, which we prefer over LAIE as it is not biased towards the later time points with larger plants.
We also test our method with different procedural models to investigate the importance of having realistic shape priors. The results can be found in the supplementary, along with a discussion of limitations.
| Net Photosynthesis | Noon (12 PM) | Late Afternooon (4 PM) |
@@ 4.3 Photosynthesis Simulation
The outputs of our method can be directly used with radiative transfer models to predict photosynthesis rates, which directly impacts crop productivity. We showcase this using Helios [3], a state-of-the-art biophysical modeling framework. Helios can predict the net photosynthesis rate for every leaf mesh face at any point in time given a set of environmental variables measured by a flux tower with sensors for temperature, humidity, radiation, etc. We show visualizations of the per-face photosynthesis rate as well as timeseries of the aggregate rate across the entire canopy in Fig. 7. Ground-truth values for the photosynthesis rate are usually calculated at a landscape-level using eddy covariance data from a flux tower [40]. Since we do not have paired flux tower measurements for our dataset (collected in 2023), we use measurements from a soybean field in the previous year (2022) with a similar location and climate.
The predictions appear to accurately represent the net photosynthesis over time as well as its spatial variations throughout the canopy, validating our pipeline’s potential for facilitating large-scale monitoring of crop productivity.
@@ 5 Conclusion
We presented a novel method for 3D modeling of agricultural crops that combines neural rendering with inverse procedural modeling to produce complete plant shapes despite heavy occlusion. Our method first constructs a neural radiance field (NeRF) and then optimizes a procedural model to be consistent with the NeRF in a visibility-aware manner. We validate our method on real-world agricultural fields and show that it can reconstruct realistic crop canopies across a variety of growth stages. We also show that it enables realistic simulations of crop photosynthesis using a radiative transfer model. The method could potentially be improved in the future by incorporating plant growth priors for temporal consistency, or optimizing the model in a coarse-to-fine manner to capture more shape details.
@@ Acknowledgments.
This work is supported by NSF Awards #2331878, #2340254, #2312102, #2414227, and #2404385, an Intel research gift, the NCSA Faculty Fellowship, and the UIUC Agroecosystem Sustainability Center.
@@
F Additional Experimental Details
F.1 Additional Procedural Model Details
@@
Our procedural soybean model is based on the mCanopy model from Song et al. [51]. Each plant consists of a main stem with up to 14 trifoliate nodes, and up to 6 branch stems with up to 2 nodes each. Unifoliate nodes and cotyldeons are ignored for simplicity. Each node consists of three coplanar leaves, with the middle leaf connected by a small petiole, and is connected to its stem by a larger petiole. The leaf length (and width), the petiole length, the angle with the stem, and the distance between nodes are variable, as they are affected by the multiplier parameters. On the main stem, when all the multipliers are set to 1, the variables follow a profile from top to bottom based on field data measured by Song et al. [51]. Branches start growing with every new main stem node starting from the 8th, with one branch per main stem node starting from the bottom and up to the 6th. The number of nodes per branch is selected randomly according to the distribution defined by Song et al. [51]. The variables for the branch nodes are the same across nodes (but are affected by the multipliers). When the number of nodes is not an integer, it is rounded up and the last node is scaled down in size linearly according to the remainder. Additionally, random noise is added to each petiole angle following a Gaussian distribution with mean 0 and standard deviation 5 degrees. Each node grows on alternating sides of their stem, with additional azimuthal rotation following a Gaussian distribution with mean 0 and standard deviation 60 degrees. The range of possible values for each model parameter are provided in Tab 3.
@@
F.2 Additional Row Fitting and Depth Rendering Details
@@
F.3 Mask and Histogram Details
@@
F.4 Baseline Implementation Details
@@
G Additional Results
G.1 Additional Qualitative Results
@@
G.2 Synthetic Scenes Evaluation
We perform additional experiments using synthetic scenes to evaluate additional aspects of canopy structure for which it is costly to obtain ground-truth in the real world. We generated 30 synthetic scenes for both soybean and maize by uniformly randomly sampling procedural model parameters and evaluated our method and the Trust-Region baseline on these scenes. We introduce two metrics to measure stem structure and topology: Node Distance Percent Error (NDPE) is the percent error on the average stem segment length between nodes (branching points), and Node Count Percent Error (NCPE) is the percent error on the total number of nodes. The results (Tab. 5) show that our method can capture these structural variables fairly accurately (but not perfectly). Note that the Poisson baseline cannot be applied here as it has no stem structure, and the MLP will simply overfit.
| Soybean | Maize | |||||
| LAIE | NDPE | NCPE | LAIE | NDPE | NCPE | |
| T-R | 2.49 | 0.48 | 0.91 | 1.46 | 0.13 | 0.70 |
| Ours | 0.97 | 0.21 | 0.23 | 0.13 | 0.05 | 0.08 |
G.3 Procedural Model Comparison
We test the performance of our method with different, simpler procedural models to investigate the importance of the procedural model’s design. We implemented the Spherical Crowns and Conical Crowns models from Helios [3] and fit them using our inverse modeling pipeline. The results (Tab. 6) show that while these models can roughly model the visible surfaces, they vastly overestimate the LAI by assuming the space underneath is densely filled with leaves. This confirms the need for grounded morphology models that capture the right assumptions about leaf arrangement.
| Soybean | Maize | |||
| LAIE | LAIPE | LAIE | LAIPE | |
| S-C | 17.85 | 1.42 | 52.12 | 6.74 |
| C-C | 53.03 | 6.13 | 46.96 | 5.91 |
| Ours | 0.69 | 0.15 | 0.97 | 0.26 |
G.4 Bayesian Optimization Performance Analysis
@@ G.5 Wind Simulation
H Limitations
One limitation of our method is that the inverse procedural modeling is dependent on the NeRF reconstruction performance, which may degrade due to wind-induced leaf motion. Another limitation is that the procedural generation models used cannot model certain details, e.g. damaged leaves and non-leaf organs of the plants. A final limitation is that our RANSAC-based row-fitting method is sensitive to hyperparameters such as the inlier threshold and segmentation threshold, which may need to be tuned per dataset. Improving the robustness by leveraging learning-based methods is a potential direction for future work.
![[Uncaptioned image]](x6.png)
![[Uncaptioned image]](x7.png)
![[Uncaptioned image]](x8.png)
![[Uncaptioned image]](fig/photosynthesis/noon_0627.png)
![[Uncaptioned image]](fig/photosynthesis/aft_0627.png)
![[Uncaptioned image]](x9.png)
![[Uncaptioned image]](fig/photosynthesis/noon_0711.png)
![[Uncaptioned image]](fig/photosynthesis/aft_0711.png)
![[Uncaptioned image]](x10.png)
![[Uncaptioned image]](fig/photosynthesis/noon_0720.png)
![[Uncaptioned image]](fig/photosynthesis/aft_0720.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230627_S1_poisson.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230627_S1_learning_baseline_final.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230627_S1_trb.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230627_S1_sb4_n125.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230711_S1_poisson.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230711_S1_learning_baseline_final.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230711_S1_trb.png)
![[Uncaptioned image]](fig/supp_qual_comparison/20230711_S1_sb4_n125.png)